# Theory of Mind abilities of Large Language Models in Human-Robot Interaction : An Illusion?

### 10 Jan 2024 v1

@@ D.4. USAR Domain
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There are two med-kits: a) medkit1 - heavier & lies closer to the room where CommX is, and b) medkit2 - lighter & lies across the hallway close to the room where a patient is located. The observer (you) has the top-view of this setting, and do not know about the properties of the med-kits. Goal : Agent has to pickup a med-kit and hand it over to CommX in the shortest plan possible. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot picks up medkit2 and hands it over to CommX. The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent should have picked up medkit1 and handed it over to CommX. 2) Because the agent should pick up the medkit which is farther away. 3) Because no plan is explicable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the medkit and turns LEFT. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent is going to the patient room directly. 2) Because the agent is going to CommX to handover the medkit. 3) Because CommX can not take medkit to the patient room.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will go to the patient room from the LEFT path. 2) Because the robot will go to the patient room from the RIGHT path. 3) Because all paths are predictable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the medkit and is now at an equal distance from the patient room, and CommX. The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can possibly go to either the patient room, or to CommX. 2) Because the robot will only go to the patient room. 3) Because the robot will only go to CommX.

### 17 Jan 2024 v2

@@ D.4. USAR Domain
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There are two med-kits: a) medkit1 - heavier & lies closer to the room where CommX is, and b) medkit2 - lighter & lies across the hallway close to the room where a patient is located. The observer (you) has the top-view of this setting, and do not know about the properties of the med-kits. Goal : Agent has to pickup a med-kit and hand it over to CommX in the shortest plan possible. The human observer can not see the robot acting in the environment. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot picks up medkit2 and hands it over to CommX. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. The human observer can not see the robot acting in the environment. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the medkit and turns LEFT. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. The human observer can not see the robot acting in the environment. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable?Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. The human observer can not see the robot acting in the environment. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the medkit and is now at an equal distance from the patient room, and CommX. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.